Suppose you have a 66.0 nC point charge and you measure a potential of 291 V. At
what distance from the charge did you make the measurement?
1
Expert's answer
2020-05-08T16:12:52-0400

The potential of the point charge is:

"\varphi = k\dfrac{q}{r}" ,

where "q" is the charge, "r" is the distance from charge to the point of measurement and "k = 9\u00d710^9 N\u22c5m^2\u22c5C^{\u22122}" is the Coulomb's constant. Express "r" from this formula ans substitute numerical values:

"r = k\dfrac{q}{\varphi} = 9\cdot 10^9 \dfrac{66\cdot 10^{-9}}{291} = 2.04 m".


Answer. r = 2.04 m.